Circular No. 6005

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994S IN NGC 4495 B. Skiff, Lowell Observatory, reports the visual discovery on June 4

UT by Larry Mitchell, Houston, TX, of a supernova (mag 14.5) located 30′′ west of the center of NGC 4495 (α = 12^{h}28^{m}.8, δ = +29^{o}26′, equinox 1950.0). A June 9 CCD observation by Skiff with the Lowell 1.1-m Hall telescope reveals the new object (V ∼ 14) ∼ 20′′ from the galaxy’s center in p.a. 250^{o}, situated along an outlying spiral arm of fairly low surface brightness. Skiff notes that, on Palomar sky survey photographs, there is also a red star of mag B ∼ 17–18 (R ∼ 15–16) located 30′′ from the galaxy’s center toward p.a. 290^{o}.

J. C. Wheeler, D. Wills, and A. Clocchiatti, University of Texas, report: “A reduced spectrum (range 507–787 nm, resolution 1.2 nm, uncalibrated in flux), obtained on June 9.24 UT with the McDonald 2.7-m telescope, shows the object (offset 13′′ west and 7′′ south from the galaxy’s center) to be a type-Ia supernova close to maximum light. The expansion velocity determined from the Si II and S II absorptions is ∼ 11 000 km/s. The redshift of the galaxy measured from Hα emission is z = 0.0152. Narrow Na D lines in absorption at this redshift (EW = 0.123 nm) suggest the foreground absorption is at least 0.8 mag in V.”

SUPERNOVA 1993J IN NGC 3031 Clocchiatti and Wheeler also communicate: “Spectra of SN 1993J from

McDonald Observatory show that emission in Hα is increasing. By May 4, the flux in Hα was greater than in the [O I] or Ca II] doublets. By June 4, the peak intensity in Hα was larger than for Ca II] at 729.1 and 732.4 nm, and, pending calibration through photometry, the total flux in Hα was larger than for May 4—implying that the supernova could be brightening in R_{c}. The line can be approximated by a flat-topped profile, centered at 656.4 nm, with FWHM = 36.5 nm. The observations suggest that Hα forms in an expanding shell heated by a mechanism other than radioactive decay, probably interaction with circumstellar matter. The presence of Hα at late times is a distinctive feature that differentiates SN 1993J from the prior transition event SN 1987K. Observations over all wavelengths before M81 becomes inaccessible are urged.”

1994 June 9 Daniel W. E. Green